Grade 12th passMechanics

the resultant of two forces one of which is double the other is 260N .if the direction of the larger force is reversed and other remains unaltered, the resultant reduces to 180N.determine the magnitude of forces and angle between the forces

Sirisha Thota
9 Years agoGrade 12th pass

2 Answers

Dhruv Arora
9 Years ago
Let the forces be of magnitude x and 2x. Then, 
x+4x2+4x2cosy=(260)2
and similarly 
x2+4x2-4x2cosy=(180)2
Simplifying we get,
10x2=100000 or x=100
and 
8x2cosy=35200
or cosy=44/100
so forces are 100N and 200N and at an angle of somewhere round 64 degrees.
vishwanath mathapati
4 Years ago
the resultant of 2 forces one of which its double the other is 260n if the direction of larger force is reversed and the other remains unattracted ,reversed reduces to the determine magnitude of force and angle between forces